Susan is 34 years old and her daughter 10 years. After how many years will she be twice as old as her daughter?

Found 2 solutions by askmemath, bonster:
Answer by askmemath(368)

Let Susan's age double her daughter's age after X years
Then
34+X = 2(10+X)
34+X = 20+2X
2X-X = 34-20
X= 14
i.e. after 14 year's Susan's age will be double her daughter's age
Check:
Susan's age = 34
After 14 year's = 34+14 = 48
Daughter's age = 10
After 14 year's = 10+14 = 24
Double 24 = 24x2 =48 = Susan's age Yay!!!

Answer by bonster(299)

10:34
11:35
12:36
13:37
14:38
15:39
16:40
17:41
18:42
19:43
20:44
21:45
22:46
23:47
24:48

48-34=14
In 14 years she will be twice as old as her daughter